OpenStudy (anonymous):

Find the solution for sin^2x = -2 + 3 cosx I did this: 1 - cos^2 = -2 + 3cosx -cos^2x = -3 + 3 cosx cos^2x + 3 cosx -3 (x )(x )

OpenStudy (anonymous):

now let cosx = t, then t^2+3t-3=0 solve this. Later come back to x

OpenStudy (phi):

cos^2x + 3 cosx -3 should be cos^2x + 3 cosx -3 = 0

OpenStudy (anonymous):

yes but that doesn't factor

OpenStudy (anonymous):

I got It!

OpenStudy (aum):

Use the quadratic formula to solve.

OpenStudy (anonymous):

i did quadratic formula

OpenStudy (anonymous):

What about cosx^2= 3+4sinx i got it down to sinx^2+4sinx+2 did quadratic and got -.626 and 3.767 but my answer is incorrect for some reason